# Finite temperature dopant-induced spin reorganization explored via tensor networks in the two-dimensional $t$-$J$ model

### v2

@@ Methods
Neighborhood tensor update (NTU)
@@
Multiplet corner transfer matrix renormalization group (mCTMRG)
@@
Data Availability
The data used in the main text of this manuscript can be found in Ref. [73]. The data used in the appendices can be given upon reasonable request.
Acknowledgements.
Figure 8: Chemical potential $\mu$ versus temperature $T$.
Appendix A Chemical Potential
For a fixed density $n$, the chemical potential is plotted versus temperature $T=1/\beta$ in Fig. 8. Similar to Refs. [23, 47], we find that the chemical potential has a linear dependence on temperature for high temperature and high filling,
@@ Appendix A Chemical Potential
Appendix B Comparison between the magnetic susceptibility and specific heat of iPEPS and METTS
@@
Appendix C Additional correlators
Here, we demonstrate an example of calculating a 5P correlator. We first generalize the $B_{ab}$} introduced in Eq. (6) to $\mathcal{B}_{abcd}\equiv S_{a}^{z}S_{b}^{z}S_{c}^{z}S_{d}^{z}$. Then, following a similar logic as in Eq. (6), we define a 5P correlator as shown in the inset of Fig. 11. The spin operators are sitting at the 4 nearest-neighbor sites of the hole operator. For simplicity, we denote this particular kind of conditional correlator as $C^{(5)}$. Fig. 11, shows $C^{(5)}$ at $T=1/2,1/6,1/10$.
Appendix D Charge clustering at intermediate temperatures in the ttJJ model via METTS
Figure 13: Cluster–size histograms from METTS snapshots resolved by hole mass – Stacked bars of the density-weighted probabilities $p_{m}$ versus cluster size $m$ for the $24\times 4$ $t$-$J$ cylinder at $T\in\{2.000,\,0.250,\,0.125,\,0.100\}$. Colors encode the cluster hole mass $\rho(\mathcal{C})=\sum_{r\in\mathcal{C}}n_{h}(r)$ binned into the non-overlapping windows $[0,0.8)$, $[0.8,1.8)$, $[1.8,2.8)$, $[2.8,3.8)$, $[3.8,4.8)$, $[4.8,5.8)$; the total height at each $m$ equals $p_{m}$. Nearest-neighbor connectivity is used with periodic wrap across the width (cylinder) and open boundaries along the length.
@@ Appendix D Charge clustering at intermediate temperatures in the ttJJ model via METTS
Fig. 13 presents the stacked histograms of $p_{m}$ versus $m$ at T=2.000, 0.250, 0.125, 0.100T=2.000,\,0.250,\,0.125,\,0.100, where the color of each bar segment encodes the window $I\in\mathcal{I}$ and the total bar height equals $p_{m}$. These windows are asymmetric intervals of width $1.0$ chosen so that the observed lobe maxima of $p_{m}$ vs. $m$ fall roughly near the middle of the corresponding window that labels their dominant total hole mass. Importantly, this coloring does not modify $p_{m}$ itself. At $T=2.000$ the distribution is nearly exponential and dominated by $m=1$ with $\rho(\mathcal{C})\in[0.8,1.8)$. Upon cooling, the weight shifts to larger $m$ and broad lobes emerge. The mass-resolved stacks show that the successive lobes of $p_{m}$ draw most of their weight from successive integer-anchored windows $I_{k}$, i.e., the clusters tend to accumulate approximately one additional hole as $m$ increases from lobe to lobe. In other words it is the cluster mass that is near-integer on a per lobe basis, not an enforced binning artifact. This pattern is compatible with an AFM background that penalizes extended domain walls, so aggregating carries step wise (one by one) lowers the exchange cost. The evolution of $p_{m}$ mirror the idea of forestalled phase separation crossover seen in the Hubbard model [28]: clusters proliferate and grow but remain finite at intermediate temperature regimes above the onset of stripes.

### v3

@@ Appendix A Numerical Details
A.1 Neighborhood tensor update (NTU)
@@
A.2 Multiplet corner transfer matrix renormalization group (mCTMRG)
@@
We used $\chi=80,100,120$ for $D=25$ and checked the energy difference between the last two iterations. The energies converge at the level of $10^{-6}\sim 10^{-8}$ even at low temperatures. After considering the multiplet structure, the final bond dimension $\chi_{f}$ can be $50\%$ larger than $\chi$.
Appendix B Chemical Potential
Figure 8: Comparison of magnetic susceptibility between different biased fields – Red lines and the blue line shows the results of biased field $h_{z}=0.1$ and $h_{z}$=0.01 at doping $n=0.9375$ respectively.
For a fixed density $n$, the chemical potential is plotted versus temperature $T=1/\beta$ in Fig. LABEL:fig:T_vs_mu. Similar to Refs. [30, 63], we find that the chemical potential has a linear dependence on temperature for high temperature and high filling,
@@ Appendix B Chemical Potential
Appendix C Comparison between the magnetic susceptibility of different biased fields
As mentioned in Sec. III, we used two different biased fields $h_{z}$ to calculate the magnetic susceptibility. In Fig. 8, we show the magnetic susceptibility using $h_{z}=0.1$ and $h_{z}=0.01$. The results agree with each other quantitatively. Note that the weaker field may be deeper in the linear regime at the cost of being more vulnerable to numerical noise.
Appendix D Comparison between the magnetic susceptibility and specific heat of iPEPS and METTS
@@
Appendix E Charge clustering at intermediate temperatures in the ttJJ model via METTS
Figure 11: Cluster–size histograms from METTS snapshots resolved by hole mass – Stacked bars of the density-weighted probabilities $p_{m}$ versus cluster size $m$ for the $24\times 4$ $t$-$J$ cylinder at $T\in\{2.000,\,0.250,\,0.125,\,0.100\}$. Colors encode the cluster hole mass $\rho(\mathcal{C})=\sum_{r\in\mathcal{C}}n_{h}(r)$ binned into the non-overlapping windows $\mathcal{I}=[0,0.8)$, $[0.8,1.8)$, $[1.8,2.8)$, $[2.8,3.8)$, $[3.8,4.8)$, $[4.8,5.8)$; the total height at each $m$ equals $p_{m}$. Nearest-neighbor connectivity is used with periodic wrap across the width (cylinder) and open boundaries along the length.
@@ Appendix E Charge clustering at intermediate temperatures in the ttJJ model via METTS
Fig. 11 presents the stacked histograms of $p_{m}$ versus $m$ at T=2.000, 0.250, 0.125, 0.100T=2.000,\,0.250,\,0.125,\,0.100, where the color of each bar segment encodes the window $I\in\mathcal{I}$ and the total bar height equals $p_{m}$. These windows are asymmetric intervals of width $1.0$ chosen so that the observed lobe maxima of $p_{m}$ vs. $m$ fall roughly near the middle of the corresponding window that labels their dominant total hole mass. Importantly, this coloring does not modify $p_{m}$ itself. At $T=2.000$ the distribution is nearly exponential and dominated by $m=1$ with $\rho(\mathcal{C})\in[0.8,1.8)$. Upon cooling, the weight shifts to larger $m$ and broad lobes emerge. The mass-resolved stacks show that the successive lobes of $p_{m}$ draw most of their weight from successive integer-anchored windows $I_{k}$, i.e., the clusters tend to accumulate approximately one additional hole as $m$ increases from lobe to lobe. In other words it is the cluster mass that is near-integer on a per lobe basis, not an enforced binning artifact. This pattern is compatible with an AFM background that penalizes extended domain walls, so aggregating carries step wise (one by one) lowers the exchange cost. Note that this lobe-like feature exists mainly at low dopings when the background antiferromagnetism is strong. The evolution of $p_{m}$ mirror the idea of forestalled phase separation crossover seen in the Hubbard model [68]: clusters proliferate and grow but remain finite at intermediate temperature regimes above the onset of stripes.
Appendix F Additional correlators
Here, we demonstrate an example of calculating a 5P correlator. We first generalize the $B_{ab}$} introduced in Eq. (6) from the main text to $\mathcal{B}_{abcd}\equiv S_{a}^{z}S_{b}^{z}S_{c}^{z}S_{d}^{z}$. Then, following a similar logic as in Eq. (6) from the main text, we define a 5P correlator as shown in the inset of Fig. 13. The spin operators are sitting at the 4 nearest-neighbor sites of the hole operator. For simplicity, we denote this particular kind of conditional correlator as $C^{(5)}$. Fig. 13, shows $C^{(5)}$ at $T=1/2,1/6,1/10$.